### v1

# LLM-Neo: Parameter Efficient Knowledge

Abstract
In this paper, we propose a novel LLM-Neo framework that efficiently transfers knowledge from a large language model (LLM) teacher to a compact student. Initially, we revisit the knowledge distillation (KD) and low-rank adaption (LoRA), and argue that they share the same paradigm. Inspired by this observation, we explore the strategy that combines LoRA and KD to enhance the efficiency of knowledge transfer. We first summarize some guidelines for this design and further develop the LLM-Neo. Experimental results on compressing Llama 2 and Llama 3 show that LLM-Neo outperforms various baselines. Further analysis demonstrates the robustness of the proposed LLM-Neo on variants of LoRA. The trained models have been available at this repository.
Index Terms:
Large Language Models, Knowledge Distillation, Low-Rank Adaption
I Introduction
Knowledge distillation (KD) [1] for large language models (LLMs)[2] is a predominant method for model compression. The key insight is to train a compact student model by mimicking the behaviors of the teacher model. One mainstreaming way is to align the logits  [3], and thus transfer the knowledge from the teacher model to the student model.
Parameter-Efficient Fine-Tuning (PEFT) [4, 5] is another commonly used technique for LLM efficiency [6]. Among various PEFT methods, the Low-rank adapter (LoRA) [7] has gained increasing popularity since it does not introduce any additional parameters for inference. During training, LoRA updates a mergeable low-rank branch instead of updating the original full parameters. Therefore, LoRA can efficiently transfer the knowledge contained in the training examples to the trained models.
Figure 1: Illustration of different knowledge transfer pipelines (KD, LoRA, and LLM-Neo). The proposed LLM-Neo pipeline combines the benefits of both the KD and LoRA approaches, that is, distilling knowledge from the teacher and low-rank branch efficiency.
In this paper, we argue that KD and LoRA follow the same paradigm, i.e., aiming at transferring knowledge while the sources differ [8]. Moreover, LoRA transfers the knowledge efficiently via the low-rank branch, while KD methods update the full parameters and typically cost much more resources. We thus ask: can we combine KD and LoRA to improve the efficiency of knowledge transfer from the teacher model?
To this end, we propose a novel LLM-Neo framework which integrates LoRA into KD to achieve parameter-efficient knowledge distillation. Specifically, as shown in Fig. 1, we follow the idea of LoRA to introduce a low-rank branch in the student model, aiming to inherit the knowledge from the teacher model. We first perform comprehensive analysis and derive valuable guidelines for the design of the proposed LLM-Neo. Experimental results on compressing Llama 2 [9] and Llama 3.1 [10] demonstrate the effectiveness and efficiency of LLM-Neo. Moreover, further analysis shows the robustness of LLM-Neo towards LoRA variants [11]. Our contributions can be concluded as follows:
  • We propose LLM-Neo, a novel approach that integrates knowledge distillation with low-rank adapter to lightweight LLMs.
  • We summarize the guidelines for LLM-Neo including: i) Typically, the larger the rank, the better, and 128 works well. ii) A learning rate close to 2e-4 works well for LoRA. iii) A larger rank requires a lower learning rate. As shown in Fig 2, trying diagonal balance is worth considering.
  • We perform extensive experiments on the Llama 2 and Llama 3.1 models to demonstrate the effectiveness of LLM-Neo. Moreover, further analyses confirm its compatibility with existing LLM optimization methods, highlighting its contributions to improved model performance and scalability.
  • We release the weights of the Llama-3.1-Neo-1B model, a small model trained on the 1 million BAAI dialogue dataset using the LLM-Neo method, contributing to the broader research and development community.
II Preliminary and Methodology
TABLE I: Comparison of SFT, LoRA, KD, and LLM-Neo on 5 benchmarks. The left represents results from Llama3.1-8B to Llama3 pruned 1B. The right for Llama2-7B to TinyLlama-1.1B. LLM-Neo achieves best average performance, with superior memory and time efficiency compared to the KD method.
Metric Llama 3.1 8B $\longrightarrow$ Llama 3 pruned 1B Llama 2 7B $\longrightarrow$ TinyLlama 1.1B
Teacher Student SFT LoRA KD LLM-Neo Teacher Student SFT LoRA KD LLM-Neo
Mem - - 63G 68G 231G 177G - - 66G 42G 167G 136G
Time - - 10min 7min 25min 20min - - 13min 12min 26min 25min
ARC-e 81.90 28.07 30.39 32.95 34.85 34.89 76.73 60.27 61.62 61.53 61.20 60.52
CEVAL 53.94 25.33 25.63 24.15 23.63 24.00 34.47 24.96 21.17 23.85 23.70 25.11
HellaS. 59.10 26.00 26.67 26.67 27.08 27.14 56.47 44.99 46.07 45.66 45.89 45.53
PIQA 80.09 53.92 54.41 56.09 57.45 56.58 78.35 74.34 72.69 72.58 73.18 72.31
WinoG. 73.72 50.43 51.38 51.85 52.64 52.64 71.03 58.72 59.27 59.91 59.67 60.54
Avg. 69.35 36.35 37.58 38.34 39.13 39.21 63.41 52.66 52.16 52.71 52.73 52.80
In this section, we first mathematically analyze the knowledge distillation (KD) and low-rank adapter (LoRA). We then extend the KD process by applying the LoRA gradient to the KL divergence term, and thus propose the LLM-Neo framework.
II-A Knowledge Distillation
In KD, the student model $S$ learns from both the ground truth labels $Y_{\text{dataset}}$ and the teacher model’s predictions $T_{\text{output}}$. For the same input, both $S$ model and $T$ model predict the next-token, producing probability distributions. The KD loss combines two parts: cross-entropy loss with the ground truth and Kullback-Leibler (KL) divergence to align the student’s predictions with the teacher’s:
$\mathcal{L}_{KD}=\alpha\cdot\mathcal{L}_{CE}(S,Y_{\text{dataset}})+(1-\alpha)\cdot\mathcal{L}_{KL}(S,T_{\text{output}}),$ (1)
where $\alpha$ is the loss weight. In this setup, the model parameters $W$ are typically updated via the gradients:
$W\leftarrow W-\eta\cdot\frac{\partial\mathcal{L}_{KD}}{\partial W}.$ (2)
II-B Low-Rank Adaption
LoRA introduces a low-rank branch to model the weight updates $\Delta W$. The update is formulated as:
$W^{\prime}=W+\Delta W,$ (3)
where $\Delta W=AB^{T}$, with $A\in\mathbb{R}^{d_{1}\times r}$ and $B\in\mathbb{R}^{d_{2}\times r}$, and $r\ll d$. Therefore, the updated parameters reduce from $d_{1}\times d_{2}$ to $(d_{1}+d_{2})\times r$. In the Supervised Fine-Tuning (SFT) process, the model learns from the training samples by minimizing the cross-entropy loss $\mathcal{L}_{SFT}$, as defined below:
$\mathcal{L}_{SFT}=\mathcal{L}_{CE}(S,Y_{\text{dataset}}).$ (4)
Hence, we can update the low-rank branch $\Delta W$ via:
$\Delta W\leftarrow\Delta W-\eta\cdot\frac{\partial\mathcal{L}_{SFT}}{\partial\Delta W}.$ (5)
II-C Extending LoRA Gradients to KL Loss
Comparing the definitions, we can conclude that KD and LoRA follow the same paradigm transferring knowledge. In Equation 1, the first term corresponds to the supervised fine-tuning loss in Equation (4), where both update the model using knowledge derived from the ground truth labels. However, the second term in the KD loss, $\mathcal{L}_{KL}(S,T_{\text{output}})$, originates from the teacher model’s outputs, representing a different source of knowledge. Therefore, it is feasible to extend the low-rank updating to the KL term as follows:
$\Delta W\leftarrow\Delta W-\eta\cdot\frac{\partial\mathcal{L}_{KL}(S,T_{\text{output}})}{\partial\Delta W}.$ (6)
This ensures that both the cross-entropy loss and the KL divergence term are optimized using LoRA’s low-rank parameters, while maintaining the efficiency of knowledge transfer through different knowledge sources.
LLM-Neo: Combining LoRA and KD
Finally, we unify the LoRA and the KD to form the LLM-Neo method. The final loss is a weighted combination of the cross-entropy loss from the dataset and the KL divergence between the student and teacher models. Specifically, the optimization process for LLM-Neo is:
$\frac{\partial\mathcal{L}_{KD}}{\partial\Delta W}=\frac{\partial\mathcal{L}_{CE}(S,Y_{\text{dataset}})}{\partial\Delta W}+\frac{\partial\mathcal{L}_{KL}(S,T_{\text{output}})}{\partial\Delta W},$ (7)
$\Delta W\leftarrow\Delta W-\eta\cdot\frac{\partial\mathcal{L}_{KD}}{\partial\Delta W}.$ (8)
In LLM-Neo, both terms contribute to updating $\Delta W$, as expressed in equations (4) and (6). Therefore, LLM-Neo retains LoRA’s parameter efficiency by applying low-rank updates across both supervised learning and knowledge distillation, combining the benefits of dataset learning and teacher model knowledge transfer.
III Experiment
III-A Experimental Implement
For the training data, we employ the BAAI Infinity-Instruct dataset [12] and randomly sample 1,000,000 samples as fine-tuning data, which include around 500M tokens. Due to limited resources, we further randomly sampled several smaller subsets of BAAI Infinity-Instruct, specifically 10,000 and 100,000 samples, corresponding to approximately 5M and 50M tokens, respectively.
For evaluation, we employ several standard benchmarks including MMLU [13], CMMLU [14], C-Eval [15], and zero-shot accuracy on commonsense reasoning datasets (i.e., PIQA [16], HellaSwag [17], WinoGrande [18], ARC-easy [19], ARC-challenge [19], and OpenbookQA [20]) using the lm-evaluation-harness package [21].
We set the maximum length as 512 for the input and keep the batch size as 4 and the gradient accumulation steps as 16 for all experiments. It takes about an hour to train on 6 Nvidia A100 40G GPUs for 2 epochs.
Figure 2: Grid search results from Llama 2 to TinyLlama for rank (2 to 256) using the 50M token subset. The matrix shows the average of 10 evaluation metrics, with darker colors indicating better performance. The results suggest that higher ranks are worth exploring with a small learning rate.
III-B Guidelines
We first distill the Llama 2 7B [9] to the TinyLlama 1.1B [22] on the 100,000 training samples. We conduct grid search for the learning rates (1e-4, 2e-4, 5e-4 and 1e-3) and rank $r$ (2, 4, 6, 8, 16, 32, 48, 128, and 256). Fig. 2 shows the average evaluation results on all the benchmarks. Based on the results, we can conclude the following guidelines:
  • Typically, the larger the rank, the better. Different with fine-tuning, KD shows a significant trends that larger rank would bring better results.
  • A learning rate close to 2e-4 works well for LoRA, and the larger one (such as 1e-3) would lead to performance decrease.
  • Considering the relationship between rank and learning rate, a larger rank requires a lower learning rate.
III-C Main Results
Based on the guidelines, we set the rank as 128 and learning rate as 2e-4. For more settings, we follow state-of-the-art LLM pruning methods [23, 24] to create our own 1B-sized student model based on the Llama 3.1, denoted as Llama 3 pruned 1B. For Llama2 series experiments, we employ Llama2-7B-instruct as the teacher models and TinyLlama-1.1B as the student model. For Llama3.1 series experiments, we employ Llama3.1-8B-instruct as the teacher and the Llama 3 pruned 1B as the student model. The minimum learning rate is set to 1e-5 for LoRA-base experiments and 1e-6 for SFT and KD experiments. We employ the 50M tokens dataset to perform the distillation.
As shown in Table I, the results of the comparative analysis of SFT, LoRA, KD, and LLM-Neo demonstrate that our LLM-Neo approach outperforms KD in memory efficiency and training time, while also surpassing SFT and standard LoRA methods in overall performance. Specifically, LLM-Neo gets an average score of 39.21 regarding the Llama 3.1, which is 0.87 score higher then the LoRA and 0.08 higher than the KD. Considering the Llama 2, LLM-Neo also outperform the baselines. Meanwhile, LLM-Neo can save about 25% GPU memory and training time, demonstrating the efficiency of knowledge transferring. Experiments on the Llama2 and Llama3.1 series further prove the robustness of our proposed LLM-Neo.
IV Extensive Analysis
IV-A Strengthen with LoRA variants
We evaluate the robustness of LLM-Neo towards LoRA variants. Specifically, we adopt the latest variant of MoSLoRA [11], which improves LoRA via mixing the subspaces. As shown in Fig. 3, LLM-Neo-MoSLoRA gets better performance than vanilla LoRA consistently, which demonstrates the robustness of LLM-Neo.
Figure 3: Improved MMLU and PIQA metrics for LLM-Neo-1B over 5 epochs using MoSLoRA, compared to the original LoRA, demonstrating enhanced effectiveness and scalability.
Figure 4: Performance with more tokens range increased from $10^{6}$ to $10^{8}$. The performance would be better with more training tokens.
IV-B Scaling law in LLM-Neo
Following the Llama3 report, we also scale the dataset and apply LLM-Neo to larger datasets progressively, including 100K, 200K, 500K, and 1M training samples. As shown in Fig. 4, the results indicate a consistent improvement in performance as the data size increased, suggesting that LLM-Neo effectively leverages larger datasets and efficiently transfers knowledge, which highlights its scalability and robustness when applied to extensive training tasks.
Figure 5: Normalized performance on 10 ability dimensions for Minitron-4B-Depth-Base model finetuned on 50M tokens using the LLM-Neo method. The chart illustrates enhancements across multiple evaluation metrics, with the expanded area indicating overall performance gains.
IV-C Applied on Minitron
We further perform distillation from Llama 3.1 to the Nvidia Minitron 4B [25, 26] using the 50M tokens. All the experiments are conducted on 6 A100 40G GPUs. Fig. 5 shows the results on several benchmarks. We can find that our proposed LLM-Neo performs well on the pruned Minitron 4B, demonstrating its robustness. Specifically, LLM-Neo gets an average score of 54.37, which is 0.52 higher than the base model. The trained weight has been available at this repository.
TABLE II: Memory usage and training time comparison across methods for the Minitron-4B-Width-Base model. Results show LLM-Neo compatibility with existing efficient methods.
Metric Baseline Methods LLM-Neo
SFT LoRA KD ZeRO1 ZeRO2
Mem 219G 103G OOM 224G 212G
Time 35min 22min - 40min 37min
IV-D Compatibility with More Memory Optimizations
To test the robustness of LLM-Neo combined with other memory optimizations, we explore the compatibility with existing LLM optimization techniques when distilling the Minitron 4B. Specifically, we further conduct LLM-Neo with ZeRO-1 methods. As shown in Table II, we can find that LLM-Neo works well with ZeRO1 and ZeRO2, highlighting the relationship between different ZeRO levels and their impact on performance metrics such as memory consumption and time efficiency. Specifically, KD would be out-of-memory though we decrease the batch size to 1.
V Conclusion
In this work, we propose a novel LLM-Neo framework, aiming to efficiently transfer knowledge from a large language model (LLM) teacher to a compact student. We first revisit the knowledge distillation (KD) and low-rank adaption (LoRA), and find that they share the same paradigm. Therefore, we explore the strategy combining LoRA and KD to enhance the efficiency of knowledge transfer. We first summarize some guidelines and further develop the LLM-Neo, i.e., the larger rank around 128 is more suitable. Experimental results on compressing Llama 2 and Llama 3.1 show that LLM-Neo outperforms various baselines. For future work, we would like to explore the performance of LLM-Neo on more training samples and more LoRA variants.