### v2

# Llm-Neo: Parameter Efficient Knowledge Distillation for Large Language Models

Abstract
Knowledge distillation (KD) has been a predominant method for compressing Large Language Models (LLMs). In this paper, we first revisit KD and Low-Rank Adaption (LoRA) and demonstrate that they follow the same paradigm. Inspired by this observation, we propose a parameter-efficient knowledge distillation method, Llm-Neo, which integrates LoRA into KD to improve the efficiency of knowledge transfer. After that, we summarize some valuable guidelines for the hyperparameters in Llm-Neo. Experimental results on compressing Llama 2 and Llama 3.2 show that Llm-Neo outperforms various baselines. Further analysis demonstrates the robustness of the proposed Llm-Neo on variants of LoRA. The code and trained models are available at Github.
Llm-Neo: Parameter Efficient Knowledge Distillation
for Large Language Models
[…]
1 Introduction
Knowledge distillation (KD) Hinton et al. (2015) is a predominant method to compress large language models (LLMs) OpenAI (2023). The key insight is to train a compact student model by mimicking the behaviors of the teacher model. One mainstreaming way is to align the logits  Wu et al. (2024a), and thus transfer the knowledge from the teacher model to the student model.
Parameter-Efficient Fine-Tuning (PEFT) Lester et al. (2021); He et al. (2021) is another commonly used technique for LLM efficiency Han et al. (2024). Among various PEFT methods, the Low-Rank Adaption (LoRA) Hu et al. (2021) has gained increasing popularity since it does not introduce any additional parameters for inference. During training, LoRA updates a mergeable low-rank branch instead of updating the original full parameters. Therefore, LoRA can efficiently transfer the knowledge contained in the training examples to the trained models.
Figure 1: Illustration of different knowledge transfer pipelines (KD, LoRA, and Llm-Neo). The proposed Llm-Neo pipeline combines the benefits of both the KD and LoRA approaches, i.e., distilling knowledge from the teacher and low-rank branch efficiency.
In this paper, we first show that KD and LoRA follow the same paradigm, i.e., aiming at transferring knowledge while the sources differ Wan et al. (2023). Moreover, LoRA transfers the knowledge efficiently via the low-rank branch, while KD methods update the full parameters and typically cost much more resources. We thus ask: can we combine KD and LoRA to improve the efficiency of knowledge transfer from the teacher model?
To this end, we propose a novel Llm-Neo framework that integrates LoRA into KD to achieve parameter-efficient knowledge distillation. Specifically, as shown in Figure 1, we follow the idea of LoRA to introduce a low-rank branch in the student model, aiming to inherit the knowledge from the teacher model. We first perform comprehensive analysis and derive valuable guidelines for the design of the proposed Llm-Neo. Experimental results on compressing Llama 2 Touvron et al. (2023) and Llama 3.1 AI@Meta (2024) demonstrate the effectiveness and efficiency of Llm-Neo. Moreover, further analysis shows the robustness of Llm-Neo towards LoRA variants Wu et al. (2024b). Our contributions can be concluded as follows:
  • We show that KD and LoRA share a unified paradigm, and thus propose a novel method, Llm-Neo, to improve the efficiency of knowledge transfer from the teacher model.
  • We summarize three key guidelines for Llm-Neo, namely: i) typically a higher rank makes better accuracy while harder to converge; ii) a learning rate about 2e-4 works well; iii) a larger rank requires a lower learning rate.
  • We perform extensive experiments on the Llama 2 and Llama 3 models to demonstrate the effectiveness of Llm-Neo. Further analyses highlight the effectiveness of improving model performance and scalability.
2 Methodology
2.1 Unified Paradigm for KD and LoRA
We can decompose the model parameters $W_{t}$ at step $t$ as: $W_{t}=W_{0}+\Delta W_{t}$, where $W_{0}$ denotes the vanilla pretrained weights and $\Delta W_{t}$ for the weight difference. Thus, the unified paradigm for updating parameters ($W_{t}\rightarrow W_{t+1}$) is:
$\underbrace{W_{t+1}}_{\text{Updated}}\leftarrow{W_{t}}-\eta\cdot\frac{\partial{\mathcal{L}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{D}}}}{\partial(W_{0}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}f}(\Delta W_{t}))}.$ (1)
In this equation, two key components emerge: 1) loss function on the information source $\mathcal{D}$ and 2) mapping function $f(\cdot)$.
KD.
KD aims to transfer the knowledge from teacher to student, which can be formulated as:
$W_{t+1}\leftarrow W_{t}-\eta\cdot\frac{\partial\mathcal{L}_{KD}}{\partial(W_{0}+\Delta W_{t})},$ (2)
$\mathcal{L}_{KD}=\alpha\cdot\mathcal{L}_{CE}(y,z_{s})+(1-\alpha)\cdot\mathcal{L}_{KL}(z_{t},z_{s}),$ (3)
[…]
Method $\bm{\mathcal{D}^{\dagger}}$ $\bm{f(x)}$ Teacher’s Parameter
Guidance Efficiency
LoRA GT $x-W_{0}$
KD GT + $\mathcal{T}$ $x$
Llm-Neo GT + $\mathcal{T}$ $x-W_{0}$
Table 1: Comparison of LoRA, KD, and Llm-Neo. GT denotes the ground truth and $\mathcal{T}$ for the teacher. We highlight the unified advantages of Llm-Neo in combining teacher guidance with parameter-efficient low-rank adaptation.
LoRA.
LoRA constrains updates to low-rank matrices. The optimization process is defined as:
$W_{t+1}\leftarrow W_{t}-\eta\cdot\frac{\partial\mathcal{L}_{CE}(z_{t},y)}{\partial(W_{0}+\Delta W_{t}-W_{0})},$ (4)
[…]
Summary.
Both LoRA and KD follow the paradigm defined in Equation 1, while the data source $\mathcal{D}$ and mapping function $f(\cdot)$ differ.
2.2 Llm-Neo Method
[…]
$W_{t+1}\leftarrow W_{t}-\eta\cdot\frac{\partial\mathcal{L}_{KD}}{\partial\Delta W_{t}}.$ (5)
Table 1 indicates the comparison among vanilla KD, LoRA, and proposed Llm-Neo. By this design, Llm-Neo can enjoy both teacher guidance and parameter efficiency.
3 Experiments
Metric Llama 3.1-8B $\longrightarrow$ Llama 3.2-1B Llama 2-7B $\longrightarrow$ TinyLlama-1.1B
Teacher Student SFT LoRA KD Llm-Neo Teacher Student SFT LoRA KD Llm-Neo
Mem - - 63G 68G 231G 177G - - 66G 42G 167G 136G
Time - - 10min 7min 25min 20min - - 13min 12min 26min 25min
ARC-e 81.90 68.52 67.72 68.39 69.15 69.11 76.73 60.27 60.61 60.35 61.49 61.24
ASDiv 88.70 67.40 67.80 67.40 67.90 69.00 61.70 18.10 21.60 20.80 21.70 21.30
HellaS. 59.10 45.07 45.20 45.30 45.29 45.38 56.47 44.99 46.82 46.86 46.73 46.72
PIQA 80.09 73.88 74.70 74.43 74.97 74.65 78.35 74.34 72.69 72.52 73.34 73.23
WinoG. 73.72 59.27 61.96 60.69 60.54 61.25 71.03 58.72 60.22 59.83 59.91 60.54
Avg. 76.70 62.83 63.48 63.24 63.57 63.88 68.86 51.28 52.39 52.07 52.63 52.60
Table 2: Comparison of SFT, LoRA, KD, and Llm-Neo on 5 benchmarks. We perform KD from Llama 3.1-8B to Llama 3.2-1B, and from Llama 2-7B to TinyLlama-1.1B. Llm-Neo achieves the best average performance, with superior memory and time efficiency compared to KD.
3.1 Experimental Implement
For the training data, we employ the BAAI Infinity-Instruct dataset (BAAI, 2024) and randomly sample 10,000 samples as fine-tuning data (around 5M tokens).
For evaluation, we employ several popular reasoning benchmarks, including MMLU Huang et al. (2023), CMMLU Li et al. (2023), C-Eval Hendrycks et al. (2021), PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2019), ARC-easy Clark et al. (2018), ARC-challenge Clark et al. (2018), and OpenbookQA Mihaylov et al. (2018)) using the lm-evaluation-harness package EleutherAI (2023). Besides, we employ several benchmarks for the math ability, including ASDiv Miao et al. (2021), GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), MAWPS Koncel-Kedziorski et al. (2016), and MathQA Amini et al. (2019).
We set the maximum input length to 512, the batch size to 4, and the gradient accumulation steps to 16 for all experiments. It takes about an hour to train on 6 Nvidia A100 40G GPUs for 2 epochs.
Figure 2: Grid search results from Llama 2 to TinyLlama for rank (2 to 256). The score matrix shows the average of 10 reasoning metrics, with darker colors indicating better performance.
3.2 Guidelines
We first distill the Llama 2-7B Touvron et al. (2023) to the TinyLlama-1.1B Zhang et al. (2024) on the 100,000 training samples. We conduct a grid search for the learning rates in {1e-4, 2e-4, 5e-4, and 1e-3} and rank $r$ in {2, 4, 6, 8, 16, 32, 48, 128, and 256}. Figure 2 shows the average evaluation results on the reasoning benchmarks. We can conclude the following guidelines:
  • Typically, the larger the rank, the better. Different with fine-tuning, KD shows a significant trends that larger rank would bring better results.
  • A learning rate close to 2e-4 works well for LoRA, and the larger one (such as 1e-3) would lead to performance decrease.
  • Considering the relationship between rank and learning rate, a larger rank requires a lower learning rate.
3.3 Main Results
Based on the guidelines, we set the rank as 128 and learning rate as 2e-4. For Llama 2 series experiments, we employ Llama 2-7B-chat as the teacher model and TinyLlama-1.1B as the student model. For Llama 3 series experiments, we employ Llama 3.1-8B-Instruct as the teacher and Llama 3.2-1B-Instruct as the student model. Moreover, we also conduct knowledge distillation on Llama 3-pruned-1B, which is pruned from the Llama 3.1 8B model (Ma et al., 2023; Kim et al., 2024), and on Minitron-4B Turuvekere Sreenivas et al. (2024); Bansal et al. (2024). The minimum learning rate is set to 1e-5 for LoRA-base experiments and 1e-6 for SFT and KD experiments. We employ the 50M tokens dataset to perform the distillation.
Table 2 shows the results on Llama 3.2 1B and TinyLlama 1B. Please refer to Appendix A and B for the results on Llama 3-pruned-1B and Minitron-4B. Overall, the results demonstrate that our Llm-Neo approach outperforms KD in memory efficiency and training time, while also surpassing SFT and standard LoRA methods. Specifically, Llm-Neo gets an average score of 63.88 regarding the Llama 3.2 1B, which is 0.64 higher than the LoRA and 0.31 higher than the KD. Considering the Llama 2, Llm-Neo also outperforms the baselines. Meanwhile, Llm-Neo can save up to 25% GPU memory and training time, demonstrating the efficiency of knowledge transferring. Experiments on the Llama 2 and Llama 3 series further prove the robustness of our proposed Llm-Neo.
Figure 3: Normalized performance on 10 benchmarks for Llama 3.2-1B Instruct model before and after distillation via Llm-Neo.
More benchmarks.
Besides the 5 benchmarks shown in Table 2, we also visualize results on more benchmarks. As shown in Figure 3, we can conclude that Llm-Neo shows effectiveness and robustness on all 10 benchmarks. particularly, Llm-Neo shows strong ability on the math tasks, such as SVAMP and GSM8K datasets.
4 Extensive Analysis
We further perform extensive analysis on the setting from Llama 3.1 8B to Llama 3-pruned-1B.
4.1 Strengthen with LoRA variants
To evaluate the robustness of Llm-Neo towards LoRA variants, we try the latest variant of MoSLoRA Wu et al. (2024b), which improves LoRA via mixing the subspaces. As shown in Figure 4, Llm-Neo-MoSLoRA gets better performance than vanilla LoRA consistently, which demonstrates the robustness of Llm-Neo.
Figure 4: Comparison of vanilla Llm-Neo and Llm-Neo-MoSLoRA on MMLU and PIQA.
4.2 Data scaling law in Llm-Neo
Following the Llama3 report, we also scale the dataset and apply Llm-Neo to larger datasets progressively, including 100K, 200K, 500K, and 1M training samples. As shown in Figure 5, the results indicate a consistent improvement in performance as the data size increased, highlighting its scalability and robustness to larger datasets.
4.3 Compatibility with More Memory Optimizations
For more memory optimization strategies, we also conduct Llm-Neo on ZeRO-1 and ZeRO-2. Experimental results demonstrate its robustness towards various optimization strategies. More details can be found at Appendix C.
Figure 5: Performance with more tokens range increased from $10^{6}$ to $10^{8}$.
5 Conclusion
In this work, we propose a novel Llm-Neo framework, aiming to efficiently transfer knowledge from a large language model (LLM) teacher to a compact student. We first revisit the KD and LoRA, and provide a unified paradigm. Therefore, we explore the strategy combining LoRA and KD to enhance the efficiency of knowledge transfer. We first summarize some guidelines, i.e., the larger rank around 128 is more suitable. Experimental results on compressing Llama 2 and Llama 3.1 show that Llm-Neo outperforms various baselines. For future work, we would like to explore the performance of Llm-Neo on more tasks and models with different structures.
6 Limitations
In this study, we provide a unified framework for knowledge acquisition through KD and LoRA, and thus propose Llm-Neo as a cross-strategy that combines the advantages of both methods, which improves efficiency while maintaining capability. To evaluate Llm-Neo, we perform experiments on the Llama series and TinyLlama and report results on various benchmarks. However, there are lots of LoRA variants and we only cover MoSLoRA. Fortunately, it is easy to apply these LoRA variants on Llm-Neo framework. We leave it for future work.
7 Ethics Statement
This work aims to improve the efficiency of knowledge distillation following the idea of LoRA and verified on various LLMs. However, it also inherits the social risks of generative LLMs, such as gender and representation bias (Lucy and Bamman, 2021). Fortunately, the proposed Llm-Neo can be applied to various LLMs and we encourage deploying the risk-free LLMs to reduce the potential ethical risks.
Appendix A Results on Llama 3-pruned-1B
To demonstrate the advantages of Llm-Neo in practical applications, we employ Llama3.1-8B-Instruct as the teacher and the Llama 3-pruned-1B as the student model. The minimum learning rate is set to 1e-5 for LoRA-base experiments and 1e-6 for SFT and KD experiments. We employ the 50M tokens dataset to perform the distillation.
As shown in Table 3, the results of the comparative analysis of SFT, LoRA, KD, and Llm-Neo demonstrate that our Llm-Neo approach outperforms KD in memory efficiency and training time, which can save about 25% GPU memory and training time, while also surpassing SFT and standard LoRA methods in overall performance. Experiments on the Llama 3-pruned-1B further prove the robustness and the efficiency of knowledge transferring of our proposed Llm-Neo, and it can work well on pruned models which means Llm-Neo has orthogonality with other LLM lightweight methods.
Appendix B Results on Minitron-4B
We further perform distillation from Llama 3.1 to the Nvidia Minitron-4B Turuvekere Sreenivas et al. (2024); Bansal et al. (2024) using the 50M tokens. All the experiments are conducted on 6 A100 40G GPUs. Figure 6 shows the results on several benchmarks. We can find that our proposed Llm-Neo performs well on the pruned Minitron 4B, demonstrating its robustness. Specifically, Llm-Neo gets an average score of 54.37, which is 0.52 higher than the base model. The trained weight has been available at HuggingFace.
Figure 6: Normalized performance on 10 ability dimensions for Minitron-4B-Depth-Base model finetuned on 50M tokens using the LLM-Neo method. The chart illustrates enhancements across multiple evaluation metrics, with the expanded area indicating overall performance gains.
Appendix C Compatibility with More Memory Optimizations
To test the robustness of Llm-Neo combined with other memory optimizations, we explore the compatibility with existing LLM optimization techniques when distilling the Minitron 4B. Specifically, we further conduct Llm-Neo with ZeRO-1 methods. As shown in Table 4, we can find that Llm-Neo works well with ZeRO1 and ZeRO2, highlighting the relationship between different ZeRO levels and their impact on performance metrics such as memory consumption and time efficiency. Specifically, KD would be out-of-memory though we decrease the batch size to 1.
Metric Baseline Methods LLM-Neo
SFT LoRA KD ZeRO1 ZeRO2
Mem 219G 103G OOM 224G 212G
Time 35min 22min - 40min 37min
Table 4: Memory usage and training time comparison across methods for the Minitron-4B-Width-Base model. Results show LLM-Neo compatibility with existing efficient methods.